# Exploring and Evaluating Real-world CXL: Use Cases and System Adoption

### v1

@@ 6.1. Page Migration with Application-Level Interleaving
Evaluation setup. To study the impact of page migration, we evaluate four memory-intensive applications, including BTree (Achermann and Panwar., 2019), an in-memory index lookup; PageRank (Beamer et al., 2015) and Graph500 (Murphy et al., 2010), both graph processing applications; and Silo (Tu et al., 2013), an in-memory database engine. We run them with 64 threads. We configure the memory consumption of each application to be around 130GB. This configuration ensures a fair comparison between static page placement solutions (i.e., the NUMA first touch and interleaving): LDRAM (i.e., 50GB) is set to less than half of each application’s memory consumption, hence both solutions can fully utilize LDRAM.
@@
Metrics. We report the execution time and page migration statistics, including the number of hint faults and migrated pages. Such statistics is collected by periodically reading the Linux counters from /proc/vmstat. Those counters capture page accesses from the entire system, but primarily influenced by page migrations because of the application.
@@
Evaluation results. Page migration observation (PMO) 1. Different applications perform differently with different page migration and static page placement solutions. No single solution can get the best performance for all applications.
@@
Figure 18 shows the execution time. We observe that BTree is not sensitive to any solution (the performance variance is less than 3%), because of its highly irregular memory access patterns. PageRank achieves the best performance with the first touch without page migration, which is 88% better than any page migration solution plus the interleaving, because PageRank has a relatively small and stable set of hot pages. Graph500 achieves the best performance with Tiering-0.8 plus the interleaving, outperforming other solutions by up to 33%, because Graph500’s hot pages are scattered across memory tiers and the interleaving is helpful for improving data locality. Silo achieves the best performance with the Tiering-0.8 plus the first touch, outperforming other solutions by up to 20%. Different from Graph500 where the hot data is distributed to many pages, Silo’s hot data is gathered into fewer pages, making the first touch more effective than the interleaving.
@@
PMO 2. When using first touch, Tiering-0.8 outperforms TPP and AutoNUMA for memory tiering.
@@
Figure 18 illustrates that, using the first touch, Tiering-0.8 outperforms NO Balance, AutoNUMA, and TPP by 7%, 3%, and 31% on average, respectively. Table 4 reveals the reasons. Compared to TPP, Tiering-0.8 has 59 $\times$ fewer hint faults on average. Memory profiling in TPP can incur a large overhead, and the reduced number of hint faults in Tiering-0.8 indicates less profiling overhead, contributing to performance improvement. Compared to AutoNUMA, Tiering-0.8 has a comparable number of hint faults but exhibits large variance in the number of migrated pages. For example, in PageRank, Tiering-0.8 results in 7.6 million more migrations than AutoNUMA. This variance is due to the dynamic adjustment of the page promotion threshold in Tiering-0.8.
@@
PMO 3. The page migration based on hint faults is not integrated well with the application-level interleaving.
@@
Table 4 reveals that using page migration plus the application level interleaving, the number of hint faults is 72,721 $\times$ less than that of page migration plus the first touch on average. Our investigation reveals that when the application-level interleaving is employed, the application’s pages are placed in unmigratable regions, preventing the pages to trigger hint faults. Hence, the page migration cannot be effective.
6.2. Page Migration with Object-Level Interleaving
@@
Evaluation setup. We evaluate HPC workloads listed in Table 2. To enable fair comparison between the application- and object-level interleaving, the capacity of LDRAM is set to 40GB for FT, 100GB for MG, and 50GB for others, such that all static page placement solutions can fully utilize LDRAM for each application. The CXL memory does not have a capacity constraint, because it is the slowest memory tier. We use 32 threads on the socket 1.
@@
Figure 19 shows that without the page migration, the object-level interleaving outperforms both the first-touch and appli-
cation-level interleaving by up to 45%. However, the page migration negatively impacts the effectiveness of the object-level interleaving: when using AutoNUMA, Tiering-0.8, and TPP, the performance degrades by 46%, 88%, and 63% on average respectively, compared to using the object-level interleaving without page migration. This occurs because the object-level interleaving utilizes the additional CXL bandwidth for bandwidth-hungry objects. However, the page migration undermines this benefit.
@@
Figure 19 shows that different HPC workloads exhibit different preferences for page migration. For example, the page migration degrades performance for FT, SP and XSBench, and yields almost no performance difference in MG, regardless of using the interleaving or first touch, because these workloads have uniformly accessed working set or highly skewed and scattered hot memory region, which make hotness detection challenging. In contrast, the page migration improves the performance of BT and LU by up to 51% and 20%, respectively, since hot pages in BT and LU have good temporal and spatial locality to be detected.
@@
7. Related Work
CXL. CXL has garnered significant attention from academia and industry (Sharma et al., 2023; Li et al., 2023; Maruf et al., 2023; Sun et al., 2023; Jang et al., 2023; Yang et al., 2023a; Shan et al., 2022; Gouk et al., 2023; Zhang et al., 2023; Wahlgren et al., 2022; Sim et al., 2023; Cho et al., 2023; Li and Yao, 2023). For example, Google has explored the potential of the CXL memory in memory tiering and swapping in its cloud computing infrastructure (Levis et al., 2023), while Microsoft has developed CXL-based memory pools for public cloud platforms  (Berger et al., 2023; Li et al., 2023).
Most of the explorations of using the CXL memory leverage NUMA servers to emulate CXL memory performance (Yang et al., 2023b, a; Jang et al., 2023; Gouk et al., 2023). For example, Yang et al. (Yang et al., 2023a) overcome the memory wall with emulated CXL-enabled SSDs. There is also a surge in research focusing on specific applications of CXL memory (Berger et al., 2023; Cho et al., 2023; Lerner and Alonso, 2024; Song et al., 2023; Li and Yao, 2023; Alonso et al., 2023; Jang et al., 2023). Besides the above efforts, Sun et al. (Sun et al., 2023) demystify the CXL memory performance using genuine CXL-ready systems and devices, which is the first paper that explores the real CXL-ready systems. However, the vast of applications from other fields (such as HPC and AI) are not extensively studied, and memory tiering and its interplay with page interleaving are not studied.
@@ 7. Related Work
Tiered memory system. Different memory components can show different memory latency, bandwidth, capacity, and monetary cost. Using multiple memory components, we can build a tiered memory systems to save cost and improve memory capacity (Kommareddy et al., 2020; Ryoo et al., 2018; Yan et al., 2019; Agarwal and Wenisch, 2017; Choi et al., 2021; Lee and Eom, 2022). Using the CXL memory, it is natural to build a tiered memory where local DDR is the fast tier and the CXL memory is a slow tier. Many solutions (Maruf et al., 2022; Kumar et al., 2021; Hildebrand et al., 2020; Li and Wu, 2022; Ren et al., 2021b; Wang et al., 2019; Dulloor et al., 2016; Raybuck et al., 2021; Kannan et al., 2017; Kim et al., 2021; Weiner et al., 2022; Wu et al., 2018, 2017; Ren et al., 2021a) have been proposed to explore and leverage the tiered memory systems to improve application performance. However, the performance of those solutions on CXL-ready systems (especially their interplay with page interleaving) is not clear (Sim et al., 2023; Ryu et al., 2023; Kim et al., 2023).
8. Conclusions
@@
We characterize the performance of the real CXL memory and explore its use cases. We study the performance of the CXL memory with a spectrum of HPC applications, explore the use of CXL for LLM training and inference, and investigate how application-transparent solutions (page interleaving and memory tiering) perform with the real CXL. We also create a data object-level interleaving method that switches between the interleaving and NUMA node-preferred polices at the data object level, and demonstrate its superior performance.

### v2

@@ VI-A Page Migration with Uniform interleaving
Evaluation setup. To study the impact of page migration, we evaluate four memory-intensive applications, including BTree [2], an in-memory index lookup; PageRank [7] and Graph500 [45], both graph processing applications; and Silo [71], an in-memory database engine. We run them with 64 threads. We configure the memory consumption of each application to be around 130GB. This configuration ensures a fair comparison between static page placement solutions (i.e., the NUMA first touch and interleaving): LDRAM (i.e., 50GB) is set to less than half of each application’s memory consumption, hence both solutions can fully utilize LDRAM.
@@
Metrics. We collect the execution time and page migration statistics, including the number of hint faults and migrated pages. Such statistics is collected by periodically reading the Linux counters from /proc/vmstat. Those counters capture page accesses from the entire system, but primarily influenced by page migrations because of the application.
@@
Evaluation results. Page migration observation (PMO) 1. Different applications perform differently with different page migration and static page placement solutions. No single solution can get the best performance for all applications. The effectiveness of a solution to an application depends on the distribution of hot pages in the working set (e.g., scattered or concentrated), and variance and size of the hot page set.
@@
Figure 16 shows execution time. We observe that BTree is not sensitive to any solution (the performance variance is less than 3%), because of its irregular memory access patterns. PageRank achieves the best performance with the first touch without page migration, which is 88% better than any page migration solution plus interleaving, because PageRank has a small and stable set of hot pages. Graph500 achieves the best performance with Tiering-0.8 plus interleaving, outperforming other solutions by up to 33%, because Graph500’s hot pages are scattered across memory tiers, and interleaving is helpful for improving data locality. Silo achieves the best performance with Tiering-0.8 plus the first touch, outperforming other solutions by up to 20%. Silo implements a B-tree-like data structure gathering hot data into fewer pages, making the first touch more effective than interleaving.
@@
PMO 2. When using first touch, Tiering-0.8 outperforms TPP and AutoNUMA, because of its smaller profiling overhead and dynamic adjustment of the page promotion threshold.
@@
Figure 16 illustrates that using the first touch, Tiering-0.8 outperforms NO Balance, AutoNUMA, and TPP by 7%, 3%, and 31%. We quantify page hints faults caused by different memory tiering strategies and observe that Tiering-0.8 has 59 $\times$ fewer hint faults. Memory profiling in TPP incurs a large overhead, and the reduced number of hint faults in tiering-0.8 contributes to performance improvement. Compared to AutoNUMA, Tiering-0.8 has a comparable number of hint faults but exhibits large variance in the number of migrated pages. For example, in PageRank, Tiering-0.8 results in 7.6 million more migrations. This variance is due to adaptiveness of page promotion threshold in Tiering-0.8.
@@
PMO 3. The page migration based on hint faults is not integrated well with the uniform interleaving.
@@
We observe that using page migration plus the application level interleaving, the number of hint faults is 72,721 $\times$ less than that of page migration plus the first touch on average. Our investigation reveals that when the application-level interleaving is employed, the application’s pages are placed in unmigratable regions, preventing the pages to trigger hint faults. Hence, page migration cannot be effective.
VI-B Page Migration with Object-Level Interleaving
@@
Evaluation setup. To enable fair comparison between the uniform and object-level interleaving, the capacity of LDRAM is set to 40GB for FT, 100GB for MG, and 50GB for others, such that all static page placement solutions can fully utilize LDRAM for each application. The CXL memory does not have a capacity constraint, because it is the slowest memory tier. We use 32 threads on the socket 1.
@@
Figure 17 shows that without page migration, the object-level interleaving outperforms the first-touch and uniform interleaving by up to 45%. However, page migration negatively impacts effectiveness of the object-level interleaving: when using AutoNUMA, Tiering-0.8, and TPP, the performance degrades by 46%, 88%, and 63% on average, compared to using the object-level interleaving without page migration. This occurs because the object-level interleaving utilizes the additional CXL bandwidth for bandwidth-hungry objects. However, the page migration undermines this benefit.
@@
Figure 17 shows that different HPC workloads exhibit different preferences for page migration. For example, the page migration degrades performance for FT, SP and XSBench, and yields almost no performance difference in MG, regardless of using the interleaving or first touch, because these workloads have uniformly accessed working set or highly skewed and scattered hot memory region, which make hotness detection challenging. In contrast, the page migration improves BT and LU performance by up to 51% and 20%, respectively, since hot pages in BT and LU have good locality to be detected.
@@
VII Related Work
CXL. In 2019, Intel introduced Compute Express Link (CXL [72, 61]), an open industry-standard interconnect between processors and devices such as accelerators, memory buffers, and smart network interfaces. Since its introduction, the CXL has garnered significant attention and investment from both researchers and industry practitioners [62, 35, 44, 67, 23, 88, 60, 18, 92, 74, 64, 11, 36, 33, 65, 4, 34, 85]. For example, Google has explored the potential of CXL memory in memory tiering and swapping in its cloud computing infrastructure [34], while Microsoft has developed CXL-based memory pools for public cloud platforms [8, 35], and Meta has designed the CXL-based tiered memory system in hyperscale datacenters [44].
Emulation-based CXL study. Most of the explorations of using the CXL memory from both industry and academics leverage NUMA servers to emulate CXL memory performance [91, 88, 23, 18, 79]. For example, Yang et al. [88] overcome the memory wall with emulated CXL-enabled SSDs, Jang et al. [23] apply emulated CXL memory to the billion-scale approximate nearest neighbor search using a software-hardware co-design solution.
Performance study on genuine CXL. Recent studies analyzed the performance of actual CXL hardware [67, 68, 41]. Sun et al. [67] represent the first comprehensive effort to analyze CXL memory performance using genuine CXL-ready systems and devices. Nonetheless, a wide range of applications from some fields (such as HPC, AI inference and training) are still not extensively studied. Tang et al. [68] investigate the CXL performance in various datacenter scenarios and propose an Abstract Cost Model to estimate the cost-benefit of using CXL memory. However, their work does not address the limitations of existing memory tiering solutions. Liu et al. [41] analyze the root causes of CXL latency’s impact on system performance. They further propose a performance model designed to predict CXL-induced slowdowns, guide the uniform interleaving ratio, and regulate page migration throttling in memory tiering. Nonetheless, their performance evaluation is restricted to CPU-centric scenarios. Our work differs in multiple perspectives, as shown in Table IV: (1) workload diversity, covering both LLMs and HPC; (2) practicality of the evaluation platform, which utilizes GPUs instead of CPUs for LLM; (3) deep insights into memory tiering; and (4) performance optimization techniques, such as object-level interleaving (Sec. VI-B) and thread assignments based on bandwidth scaling (Sec. III).
@@ VII Related Work
Tiered memory systems. Different memory components [77, 83, 9, 16, 19, 1, 72, 61, 10] can show different memory latency, bandwidth, capacity, and monetary cost. Using multiple memory components, tiered memory systems can save cost and improve memory capacity [28, 58, 87, 3, 12, 31, 56]. Using the CXL memory, it is natural to build a tiered memory where local DDR is the fast tier and the CXL memory is a slow tier. Many solutions [43, 29, 20, 37, 52, 76, 17, 50, 24, 25, 78, 81, 80, 51, 75, 84, 55, 86, 40, 21, 57, 38, 39, 42, 53, 90, 89, 82] have been proposed to explore and leverage the tiered memory systems to improve application performance. However, the performance of those solutions on CXL-ready systems (especially their interplay with page interleaving) is not clear [64, 59, 26].
VIII Conclusions
@@
In this paper, we characterize the performance of the real CXL memory and explore its use cases. We explore the use of CXL for LLM training and inference, study the performance of the CXL memory with a spectrum of HPC applications, and investigate how application-transparent solutions (page interleaving and memory tiering) perform with the real CXL. We also create a data object-level interleaving method that switches between the interleaving and NUMA node-preferred polices at the data object level, and demonstrate its superior performance. We hope that our study can shed some lights on how the future HPC systems can leverage CXL memory expansion.